OpenStudy (anonymous):

find the quadratic polynomial that has a minimum value of -3 when x=-2 and passes through the point (1,6)

OpenStudy (anonymous):

vertex(-2,-3) y=a(x+2)^2-3 put (1,6) 6=a(1+2)^2 -3 9=9a a=1 therefore, the required equation is y=(x+2)^2-3

OpenStudy (anonymous):

thank you! =]